# Digital Simulation of Non-Abelian Parafermions in Superconducting Circuits

### v3

@@ VI Experimental parameters
In this section, we briefly discuss typical parameters for SC circuits to realize our scheme, particularly for the case of $d=3$, and provide an outlook on the potential applications in high-dimensional topological quantum computing. One key parameter in SC circuits is the fidelity of single- and two-qutrit gates. Recent work has demonstrated the realization of the generalized Hadamard gate for qutrits, achieving a high fidelity of $0.992$ Yurtalan et al. (2020). More challenging is the implementation of high-fidelity two-qutrit gates. In another recent study Ref. Goss et al. (2022), the authors report the realization of $CZ^{\dagger}$ and $CZ$ gates for qutrits, with fidelities of $0.973$ and $0.952$, respectively. Since the fidelity of single-qutrit gates is much higher than that of two-qutrit gates, we will, as a first approximation, neglect errors from single-qutrit gates in the following analysis. Our ground state preparation protocol requires three two-qutrit gates for each dark plaquette. To achieve a ground state preparation fidelity of $0.90$, the fidelity of the two-qutrit gates needs to be above $0.965$, which is very close to the fidelity achievable with existing technology. Achieving higher level of fidelity would require further technological advances in SC devices, highlighting an important avenue for future research in this field. Other typical parameters include the relaxation time $T^{ij}_{1}$ and the dephasing time $T^{ij}_{2}$ between different levels ($ij\in\{01,12,02\}$), which must be sufficiently long to complete the braiding, fusion procedure, and measurement. Recent studies on transmon qutrits report $T^{ij}_{1}$ and $T^{ij}_{2}$ values typically around $10\mu s$ or higher Luo et al. (2023); Goss et al. (2022). Considering that the gate times for high precision control are on the order of $10-100ns$ for single-qutrit gates and $100-500ns$ for two-qutrit gates, these parameters suggest that ground state preparation, along with subsequent fusion and braiding procedures, is highly promising and achievable with the near-future technology Luo et al. (2023).
@@ VII Discussion and conclusion
Before concluding, we discuss the potential applications in high-dimensional quantum computing. Recent developments highlight several advantages over qudit-based quantum information processing Campbell (2014) and flexible simulation of quantum dynamics Neeley et al. (2009). Qudit versions of various quantum algorithms have already been developed Nguyen and Kim (2019); Ivanov et al. (2012); Ye et al. (2011). High-dimensional quantum computing based on non-Abelian parafermions is inherently fault-tolerant due to topological protection. While realizing parafermions in physical materials remains a great challenge, our present digital simulation provides a feasible scheme for high-dimensional gate-based topological quantum computing in the near future. One exciting direction is to implement quantum algorithms in qudit systems using parafermions braidings. Another challenge is using digital simulation of parafermions to implement the topologically protected $CX$ gate, involving eight parafermions and twelve half-braidings Hutter and Loss (2016).
@@ Acknowledgments
In this section, we introduce the algebraic theory of $\mathbb{Z}_{3}$ parafermions. For a general $d$, we refer to Ref. Hutter and Loss (2016). The algebraic relations for $\mathbb{Z}_{3}$ parafermions are given by:

### v4

@@ VI Experimental parameters
In this section, we briefly discuss typical parameters for SC circuits to realize our scheme, particularly for the case of $d=3$. One key parameter in SC circuits is the fidelity of single- and two-qutrit gates. Recent work has demonstrated the realization of the generalized Hadamard gate for qutrits, achieving a high fidelity of $0.992$ [82]. More challenging is the implementation of high-fidelity two-qutrit gates. In recent study Ref. [84], the authors report the realization of $CZ^{\dagger}$ and $CZ$ gates for qutrits, with fidelities of $0.973$ and $0.952$, respectively. Since the fidelity of single-qutrit gates is much higher than that of two-qutrit gates, we will, as a first approximation, neglect errors from single-qutrit gates in the following analysis. Our ground state preparation protocol requires three two-qutrit gates for each dark plaquette. To achieve a ground state preparation fidelity of $0.90$, the fidelity of the two-qutrit gates needs to be above $0.965$, which is very close to the fidelity achievable with existing technology. Achieving higher level of fidelity would require further technological advances in SC devices, highlighting an important avenue for future research in this field. Other typical parameters include the relaxation time $T^{ij}_{1}$ and the dephasing time $T^{ij}_{2}$ between different levels ($ij\in\{01,12,02\}$), which must be sufficiently long to complete the braiding, fusion procedure, and measurement. Recent studies on transmon qutrits report $T^{ij}_{1}$ and $T^{ij}_{2}$ values typically around $10\mu s$ or higher [81, 84]. Considering that the gate times for high precision control are on the order of $10-100ns$ for single-qutrit gates and $100-800ns$ for two-qutrit gates, these parameters suggest that ground state preparation, along with subsequent fusion and braiding procedures, is highly promising and achievable with the near-future technology [81].
@@ VII Discussion and conclusion
Before concluding, we discuss the potential applications in high-dimensional quantum computing. Recent developments highlight several advantages over qudit-based quantum information processing [96] and flexible simulation of quantum dynamics with qudits [97]. Qudit versions of various quantum algorithms have already been developed [98, 99, 89]. High-dimensional quantum computing based on non-Abelian parafermions is inherently fault-tolerant due to topological protection. While realizing parafermions in physical materials remains a great challenge, our digital simulation scheme offers a promising approach to high-dimensional gate-based topological quantum computing in the near future. One exciting direction is to implement qudit quantum algorithms using parafermions braidings. Another challenge is using digital simulation of parafermions to implement the topologically protected $CX$ gate, involving eight parafermions and twelve half-braidings [56].
@@ Acknowledgments
In this section, we introduce the algebraic theory of $\mathbb{Z}_{3}$ parafermions. For a general $d$, we refer to Ref. [56]. The algebraic relations for $\mathbb{Z}_{3}$ parafermions are given by: